How to calculate the oxidation number of cobalt in K[Co(CN) (CO)2(NO)]

Grade:12

2 Answers

Arun
5 years ago
Let oxidation number of Co(cobalt) = x
Then
1 + x + (-1) + 0 +1 = 0
x = -1.
Hope it helps
Regards
Arun
Rohit
5 years ago
In K[Co(CN)(CO)2(NO)]
Let the oxidation state of Co be x.
Hence 1+x-1+0+1 =0
Hence OS of Co is -1.
Note here the OS of N is generally -3 but due to more EN O atom it becomes +3. Hence OS of NO is 3-2=+1.
Hope it helps.
Regards 
ROHIT......